# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v2

@@ 6 Bellman Optimality for the HD-S Case
$0\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{S}}}\underline{\alpha}(\mu,\pi,\kappa)-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\leq\epsilon.$
@@
The proof of Proposition 6.1 is deferred to Appendix F.1, where we instantiate the RL policy $\pi_{\mathrm{RL}}$ with the online algorithm in Zhang and Xie, [30]. This choice is illustrative rather than exclusive: any online RL algorithm that (i) uses only deterministic actions (consistent with the assumption $\{\delta_{a}:a\in A\}\subseteq\mathcal{Q}$) and (ii) achieves sublinear regret can be employed to obtain $\epsilon$-optimality for any prescribed $\epsilon>0$.
@@ Appendix D Proof of Theorem 6
Appendix E Proof of Corollary 7.3
@@
Appendix F Proofs for Section 6
F.1 Proof of Proposition 6.1
@@ Appendix F Proofs for Section 6
$0$ $\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (F.1)
@@
$0$ $\leq\alpha^{*}_{\kappa}-\overline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (F.2)
@@
Since $\pi_{\mathrm{RL}}$ only uses deterministic actions and $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, $\pi_{\mathrm{RL}}\in\Pi_{\mathrm{H}}(\mathcal{Q})$. Therefore, going back to (F.1), we have that
@@
Finally, to conclude the proposition, we note that if $\underline{\alpha}$ is replaced by $\overline{\alpha}$, the derivation in (F.1) is still valid. This, coupled with (F.2), yields the limsup version of Theorem 6.1. ∎
F.2 Proof of Theorem 8

### v3

@@ 6 Bellman Optimality for the HD-S Case
$0\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\stackrel{{\scriptstyle(ii)}}{{\leq}}\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{S}}}\underline{\alpha}(\mu,\pi,\kappa)-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)\leq\epsilon.$
@@
Note that in the second inequality, we swap to inf-sup. The proof of Proposition 6.1 is deferred to Appendix G.1, where we instantiate the RL policy $\pi_{\mathrm{RL}}$ with the online algorithm in Zhang and Xie, [39]. This choice is illustrative rather than exclusive: any online RL algorithm that (i) uses only deterministic actions (consistent with the assumption $\{\delta_{a}:a\in A\}\subseteq\mathcal{Q}$) and (ii) achieves sublinear regret can be employed to obtain $\epsilon$-optimality for any prescribed $\epsilon>0$.
@@
Below, we provide an example illustrating the case $\alpha^{*}\neq\alpha^{\prime}$. Our example is adapted from Wang et al., [32], which is visualized in Figure 1.
IGB
(a) $p^{(1)}$
IGB
(b) $p^{(2)}$
Figure 1: A robust MDP example of the case $\alpha^{*}\neq\alpha^{\prime}$, where the red line and the blue line represent actions $a_{1}$ and $a_{2}$, respectively.
The state space is $S=\left\{{\text{I},\text{G},\text{B}}\right\}$, where I stands for the initial state, G is the good state, and B is a bad state. We consider the following reward function $r$, which does not depend on the control actions
$r(\text{I})=0,\;r(\text{G})=1,\;r(\text{B})=-1.$
In I, two actions ${A}=\{a_{1},a_{2}\}$ lead to different dynamics, whereas in G and B, taking different two actions will not change the dynamics. In particular, in state I, the S-rectangular adversary can choose from $\mathcal{P}_{\text{I}}:=\left\{{p_{\text{I}}^{(1)},p_{\text{I}}^{(2)}}\right\}$. In kernel form,
$p^{(1)}(\text{B}|\text{I},a_{1})=1,\;p^{(1)}(\text{G}|\text{I},a_{2})=1,\text{ and }p^{(2)}(\text{G}|\text{I},a_{1})=1,p^{(2)}(\text{B}|\text{I},a_{2})=1.$
Here, $p^{(1)}(B|\text{I},a_{1})=1$ means that in state I, if the controller selects action $a_{1}$ and the adversary chooses $p_{\text{I}}^{(1)}$, the MDP transitions to state B with probability 1. The other transition probabilities are interpreted in the same way. We assume $\mathcal{Q}=\mathcal{P}(\mathcal{A})$.
Note that in this robust MDP instance, both the controller and the adversary are communicating. Specifically, fixing any action (or randomized policy) chosen by the controller, state G (respectively B) can reach state B (respectively G) under at least one of the kernels $p^{(i)}$, $i=1,2$, while state I is always recurrent. Similarly, fixing any adversarial kernel $p^{(i)}$, $i=1,2$, there always exists a control action that enables a transition from G (B) to B (G).
In this example, it is not hard to see that if the controller can only use stationary policies and the adversary plays second, the best stationary strategy of the controller is to randomize the two actions with probability $1/2$. Therefore, one can easily verify that $\overline{\alpha}(\text{I},\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}})=0$, and a solution to the robust Bellman equation (3.1) is
$\alpha^{*}=0,\;u^{*}(\text{I})=0,\;u^{*}(\text{G})=1,\;u^{*}(\text{B})=-1.$
In particular, Theorem 1 holds and
$0$ $=\overline{\alpha}(\text{I},\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})=\underline{\alpha}(\text{I},\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})$
$=\overline{\alpha}(\text{I},\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}})=\underline{\alpha}(\text{I},\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}})=\overline{\alpha}(\text{I},\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}})=\underline{\alpha}(\text{I},\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}).$
On the other hand, if the controller plays second, it is always able to exploit the knowledge of the adversary’s choice $p^{(i)}$, and counter with $a_{(i\text{ mod }2)+1}$. So, one would expect that $\alpha^{\prime}=0.5$, as, in the long run, the Markov chain will spend half of the time in state I and the other half of the time in G. With this intuition, it is not hard to verify that the inf-sup equation (3.2) is solved by
$\alpha^{\prime}=0.5,\;u^{\prime}(\text{I})=0,\;u^{\prime}(\text{G})=0.5,\;u^{\prime}(\text{B})=-1.5.$
Then, Theorem 8 indicates $\underline{\alpha}(\text{I},\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})=\overline{\alpha}(\text{I},\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})=0.5.$ In particular, as Corollary 8.1 suggests, stationary policies cannot be optimal for the controller for this MDP instance.
Moreover, we can construct a simple optimal history-dependent controller policy that achieves the optimal gain of $0.5$ as follows. At time 0, starting from the initial state I, the controller selects an arbitrary action, say $a_{2}$.
If state G is observed at time $1$, the controller can infer that the adversary has selected kernel $p^{(1)}$. Consequently, the controller continues to choose $a_{2}$ for all subsequent time steps. This induces a deterministic Markov chain alternating between I and G, yielding a long-run average reward of $0.5$.
Conversely, if state B is observed at time $1$, the controller infers that the adversary has selected kernel $p^{(2)}$. The controller then switches to action $a_{1}$ for all subsequent time steps, again inducing a deterministic Markov chain alternating between I and G, and thus achieving the same long-run average reward of $0.5$.
@@ Appendix D Proof of Theorem 6
Appendix E Proof of Theorem 7
Proof.
We first show statement (1).
Note that the mapping
$(\phi,p_{s})\rightarrow\sum_{a\in A,\;s^{\prime}\in S}\phi(a)p(s^{\prime}|s,a)[r(s,a)-\alpha^{\prime}+u^{\prime}(s^{\prime})]=E_{\phi,p_{s}}[r(s,A_{0})-\alpha^{\prime}+u^{\prime}(X_{1})]$
is bi-linear in $\phi$ and $p_{s}$. Since both $\mathcal{Q}$ and $\mathcal{P}_{s},\;s\in S$ are convex and one of them is compact, by Sion’s minimax Theorem (Corollary 3.3 in Sion, [26]), we have that for all $s\in S$,
$u^{\prime}(s)$ $=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}E_{\phi,p_{s}}[r(s,A_{0})-\alpha^{\prime}+u^{\prime}(X_{1})]$
$=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}[r(s,A_{0})-\alpha^{\prime}+u^{\prime}(X_{1})].$
Hence, $(u^{\prime},\alpha^{\prime})$ solves (3.1).
To prove the second statement, we define
$q^{\prime}(s,a)=r(s,a)-\alpha^{\prime}+\inf_{\nu\in\mathcal{P}_{s,a}}\sum_{s^{\prime}\in S}\nu(s^{\prime})u^{\prime}(s^{\prime}).$
Notice that since $\left\{{\delta_{a}:a\in A}\right\}\subset\mathcal{Q}$, we have
$\max_{a\in A}q^{\prime}(s,a)$ $=\sup_{\phi\in\mathcal{Q}}\sum_{a\in A}\phi(a)q^{\prime}(s,a)$ (E.1)
$=\sup_{\phi\in\mathcal{Q}}\sum_{a\in A}\phi(a)\inf_{\nu\in\mathcal{P}_{s,a}}\sum_{s^{\prime}\in S}\nu(s^{\prime})[r(s,a)-\alpha^{\prime}+u^{\prime}(s^{\prime})]$
$\stackrel{{\scriptstyle(i)}}{{=}}\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}\sum_{a\in A,\,s^{\prime}\in S}\phi(a)p(s^{\prime}|s,a)[r(s,a)-\alpha^{\prime}+u^{\prime}(s^{\prime})]$
$\leq u^{\prime}(s)$
where $(i)$ follows from SA-rectangularity that $\mathcal{P}_{s}=\bigtimes_{a\in A}\mathcal{P}_{s,a}$.
On the other hand,
$u^{\prime}(s)$ $=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}\sum_{a\in A,\,s^{\prime}\in S}\phi(a)p(s^{\prime}|s,a)[r(s,a)-\alpha^{\prime}+u^{\prime}(s^{\prime})]$ (E.2)
$=\inf_{\nu_{1}\in\mathcal{P}_{s,a_{1}}}\dots\inf_{\nu_{|A|}\in\mathcal{P}_{s,a_{|A|}}}\sup_{\phi\in\mathcal{Q}}\sum_{i=1}^{|A|}\phi(a_{i})\left[r(s,a_{i})-\alpha^{\prime}+\sum_{s^{\prime}\in S}\nu_{i}(s^{\prime})u^{\prime}(s^{\prime})\right].$
For any $\delta>0,$ choose $\nu^{\prime}_{i}\in\mathcal{P}_{s,a_{i}}$ so that
$\sum_{s^{\prime}\in S}\nu^{\prime}_{i}(s^{\prime})u^{\prime}(s^{\prime})\leq\inf_{\nu\in\mathcal{P}_{s,a_{i}}}\sum_{s^{\prime}\in S}\nu(s^{\prime})u^{\prime}(s^{\prime})+\delta.$
Then, continue from (E.2), we have that
$u^{\prime}(s)$ $\leq\sup_{\phi\in\mathcal{Q}}\sum_{i=1}^{|A|}\phi(a_{i})\left[r(s,a_{i})-\alpha^{\prime}+\sum_{s^{\prime}\in S}\nu^{\prime}_{i}(s^{\prime})u^{\prime}(s^{\prime})\right]$
$\leq\delta+\sup_{\phi\in\mathcal{Q}}\sum_{i=1}^{|A|}\phi(a_{i})\left[r(s,a_{i})-\alpha^{\prime}+\inf_{\nu\in\mathcal{P}_{s,a_{i}}}\sum_{s^{\prime}\in S}\nu(s^{\prime})u^{\prime}(s^{\prime})\right]$
$=\delta+\sup_{\phi\in\mathcal{Q}}\sum_{a\in A}\phi(a)q^{\prime}(s,a)$
$=\delta+\max_{a\in A}q^{\prime}(s,a).$
Since $\delta>0$ can be arbitrary small, combining this with (E.1), we conclude that $u^{\prime}(s)=\max_{a\in A}q^{\prime}(s,a)$ for all $s\in S$.
Finally, we note that from (E.1),
$u^{\prime}(s)=\max_{a\in A}q^{\prime}(s,a)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}\sum_{a\in A,\,s^{\prime}\in S}\phi(a)p(s^{\prime}|s,a)[r(s,a)-\alpha^{\prime}+u^{\prime}(s^{\prime})].$
Therefore, $(u^{\prime},\alpha^{\prime})$ solves (3.1), completing the proof. ∎
Appendix F Proof of Corollary 7.3
@@ Appendix E Proof of Theorem 7
Appendix G Proofs for Section 6
G.1 Proof of Proposition 6.1
@@ Appendix G Proofs for Section 6
$0$ $\leq\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}})-\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}}\underline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (G.1)
@@
$0$ $\leq\alpha^{*}_{\kappa}-\overline{\alpha}(\mu,\pi_{\mathrm{RL}},\kappa)$ (G.2)
@@
Since $\pi_{\mathrm{RL}}$ only uses deterministic actions and $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, $\pi_{\mathrm{RL}}\in\Pi_{\mathrm{H}}(\mathcal{Q})$. Therefore, going back to (G.1), we have that
@@
Finally, to conclude the proposition, we note that if $\underline{\alpha}$ is replaced by $\overline{\alpha}$, the derivation in (G.1) is still valid. This, coupled with (G.2), yields the limsup version of Theorem 6.1. ∎
G.2 Proof of Theorem 8